OpenStudy (sedatefrog712):

MEDAL!!! A right triangle has base (x - 5) units and height (4x - 10) units. Part A: What is the square of the length of the hypotenuse of the triangle? Show your work. (4 points) Part B: What is the area of the triangle? (3 points) Part C: Using the solution obtained in Part B, explain the closure property of multiplication of polynomials. (3 points)

OpenStudy (redheadangel):

Area of triangle=12×base×height= 12×(x−5)×(4x−10)=⋯

OpenStudy (redheadangel):

The closure property of polynomials says that the product of two polynomials is also a polynomial.

OpenStudy (confluxepic):

(x-5)^2 +(4x-10)^2 =hypotenuse^2

OpenStudy (sedatefrog712):

can you help me with this step by step like a comes first then b etc

OpenStudy (anonymous):

\[(x-5)^2\]

OpenStudy (sedatefrog712):

how do you get that